OpenStudy (anonymous):

find sum of the 1st 20 terms in the arithmetic series 5,8,11,14

OpenStudy (amistre64):

what do we add to get to each next term?

OpenStudy (amistre64):

called a common difference

OpenStudy (anonymous):

3

OpenStudy (amistre64):

3 is right :) so, Sn = n(first+last)/2 if i recall it right

OpenStudy (amistre64):

to determine the 20th term is a20 = 5+3(19) = 62 62+5 = 67*20/2 = 670 ... if my math was good

OpenStudy (anonymous):

Oh so it's 670? Thank you so so so much! wow thank you.

OpenStudy (amistre64):

i hope so, id dbl chk with the wolf: sum (5+3(n-1)) from 1 to 20

OpenStudy (amistre64):

670 :)

OpenStudy (anonymous):

Can you help me with one more? Find the 7th term in the geometric sequence 4,20,11

OpenStudy (amistre64):

20/4 = our common ratio an = a1 * r^(n-1) a7 = 4 * (20/5)^6

OpenStudy (amistre64):

your 11 is bad there, is that a typo?

OpenStudy (anonymous):

sorry i meant 100

OpenStudy (anonymous):

yes :/

OpenStudy (amistre64):

whew!! lol, then yeah; 4*5^6 should do it

OpenStudy (anonymous):

so a^7 = 4*5^6?

OpenStudy (anonymous):

dd you get 250?

OpenStudy (amistre64):

i think its bigger than that; 5^4 = 25^2 = 625 alone

OpenStudy (amistre64):

625*20 is what id end up with

OpenStudy (amistre64):

12500 perhaps? hard to do in the head :)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=4*%285%5E6%29 almost had it :)

OpenStudy (amistre64):

i left off the last *5 .... 12500*5=62500